A man entered an orchard through 7 gates, and there took a certain number of apples.

When he left the orchard, he gave the first guard half of the apples that he had and 1

apple more. To the guard at the second gate, he gave half of his remaining apples and 1

apple more. He did the same with each of the remaining 5 guards and left the orchard

with 1 apple. How many apples did he gather in the orchard?


1
Expert's answer
2021-12-01T14:45:52-0500

Before 7th gates a man has "(1+1)\cdot2=4" apples.

Before 6th gates a man has "(4+1)\cdot2=10" apples.

Before 5th gates a man has "(10+1)\cdot2=22" apples.

Before 4th gates a man has "(22+1)\cdot2=46" apples.

Before 3rd gates a man has "(46+1)\cdot2=94" apples.

Before 2nd gates a man has "(94+1)\cdot2=190" apples.

Before 1st gates a man has "(190+1)\cdot2=382" apples.


A man gathered 382 apples in the orchard.